Please help me solve for x in the equation y+2x=12. I am solving with solve systems by substitutions and y is equal to x.
Answer by stanbon(75887)

solve for x in the equation y+2x=12
---------------
2x = 12-y
---
x = 6 - (1/2)y
=====================
Cheers,
Stan H.
=====================